# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v3

@@ 3.2.1 The discrete Fréchet distance
Let $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$ be two sequences of points in some normed space. Conceptually, we have two agents starting at $p_{1}$ and $q_{1}$ respectively. During an atomic time interval, one of them can jump forward to the next point in the sequence (one can allow both to jump forward in the same time, but we disallow it for the sake of simplicity of exposition). In the end, both agents have to arrive to $p_{n}$ and $q_{m}$, respectively, and our purpose is to minimize the maximum distance between them during this motion.
@@
Consider two sequences $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$, a discrete morphing is an $x/y$-monotone path $\tau=(1,1),\ldots,(n,m)$ in the grid graph defined over the set of points $U=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$ from $(1,1)$ to $(n,m)$. For a vertex $(i,j)\in U$, let $\mathcalb{h}\left(i,j\right)=\left\|{p_{i}q_{j}}\right\|$ be its associated weight. The path $\tau$ is restricted to vertical and horizontal edges of the grid. The width of $\tau$ is ${\omega}_{\pi,\sigma}\left(\tau\right)=\max_{(i,j)\in V\left(\tau\right)}\mathcalb{h}\left(i,j\right).$ The discrete Fréchet distance between $\pi$ and $\sigma$ is the minimum of width of any morphing between the two sequences.
@@ 3.2.2 The retractable discrete Fréchet distance
Let $\tau=z_{1},\ldots,z_{n+m-1}\in U$ be a discrete morphing between two sequences $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$. Let $D(i,j)=\max_{k:i<k<j}\mathcalb{h}\left(z_{k}\right)$ be the inner width of the morphing $\tau(i,j)=z_{i},z_{i+1},\ldots,z_{j}$. For simplicity assume all pairwise distances $\left\|{p_{i}q_{j}}\right\|$ are distinct. The unique point $\mathcalb{b}(i,j)$ realizing the minimum $D(i,j)$, overall possible grid monotone paths $\tau^{\prime}$ between $z_{i}$ and $z_{j}$, is the bottleneck between $i$ and $j$. The retractable discrete Fréchet morphing between $\pi$ and $\sigma$ is the unique morphing $\tau$, such that
@@
Although the weights here are defined on the vertices, it is easy enough to interpret them as being on the edges (by for example, assigning a grid edge $z\rightarrow z^{\prime}$ the weight $\max(\mathcalb{h}\left(z\right),\mathcalb{h}\left(z^{\prime}\right))$. Plugging this (implicit) graph into Lemma 3.2 readily yields the following result.
@@
    The first running time bound of Lemma 3.5 is a worst case bound, and for many realistic inputs it is much smaller if the (discrete) Fréchet distance between them is small compared to the diameter of the two curves. See Figure 3.1 for an example.
@@ 4.1 Definition and basic algorithm
Let $\mathcalb{p}_{i}=\overrightarrow{p_{i}p_{i+1}}$ and $\mathcalb{q}_{j}=\overrightarrow{q_{j}q_{j+1}}$. Let $x_{i}=\left\|{\pi[p_{1},p_{i}]}\right\|$ and $y_{j}=\left\|{\sigma[q_{1},q_{j}]}\right\|$. The grid cell $C_{i,j}$ is the subrectangle $[x_{i},x_{i}+\left\|{p_{i}p_{i+1}}\right\|]\times[y_{j},y_{j}+\left\|{q_{j}q_{j+1}}\right\|]$. The elevation function for any point $(x,y)$ inside $C_{i,j}$ is given by
@@
We emphasize that that the running time bound in this theorem is pessimistic, and in many cases the algorithm is significantly faster. For an example of the output of the algorithm, see Figure 4.2.
Figure 4.2: The VE-Fréchet morphing for the two curves from Figure 1.3. Note, that the solution is “slightly” not $x/y$-monotone. For animation of this morphing, follow the link.
@@ 4.1.1 Reachable complexity
A similar notion (relative free space complexity) was used by Driemel et al. dhw-afdrc-12. If $r=\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)$, then the algorithm for computing the VE morphing only explores vertices in the $r$-reachable region, which readily implies the following.
@@
The above implies that the worst case, for the above algorithm, is when $N=nm$. This happens for example (somewhat counterintuitively) when $\pi$ and $\sigma$ are far from each other, and their VE-Fréchet distance is realized by the distance of their endpoints (say, in the start of the morphing). This somewhat absurd situation can happen for real world inputs – follow this link for an example. Furthermore, this quadratic behavior is probably unavoidable, as computing the VE-Fréchet distance probably requires quadratic time in the worst case. We offer some approaches to avoid this for real world input later on.
@@ 4.2 Basic properties of the VE Fréchet morphing/distance
Specifically, let $\mathcalb{m}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. If we project $\mathcalb{m}$ on the $x$-axis, denoted by $x(\mathcalb{m})$, we get a path from 00 to $\left\|{\pi}\right\|$ on the real line. Note, that this path is not necessarily monotone, but importantly, for any vertex $v\in\pi$ it is monotone. Specifically, let $\alpha$ be the distance of $v$ along $\pi$ from its start. Then, the path $x(\mathcalb{m})$ crosses from one side of $\alpha$ to the other side, exactly once (assuming $v$ is an internal vertex). To see why this is true, observe any morphing must cross the vertical line $x=\alpha$. This line has vertices along it, and observe that in $G_{\textsf{VE}}\left(\pi,\sigma\right)$ it is only possible to cross this line by using one of these vertices, as no edge crosses from one side of this line to the other (i.e. the vertices on this line are a separator). However, vertices along this line have outgoing edges only to vertices that are to the right (or remain on the same $x$, but these target vertices are on grid edges adjacent on to the right to this line).
@@ 5.1 Direct Monotonization
If the VE-Fréchet morphing is $x/y$-monotone, then it realizes the (regular) Fréchet distance between the curves. Indeed, if the morphing is monotone then it is a valid morphing considered by Definition 1.4 for the Fréchet distance, and thus $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)\geq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\geq\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)$, where the latter inequality is from Lemma 4.8.
@@
A VE-Fréchet morphing $\mathcalb{m}$ in a free space diagram $R=R\left(\pi,\sigma\right)$ of two curves $\pi$ and $\sigma$ with $n$ and $m$ vertices, respectively, is a polygonal path with $t=n+m+O(1)$ vertices. Thus, a morphing can be represented as $\mathcalb{m}=(x_{1},y_{1}),\ldots,(x_{t},y_{t})$. A consecutive pair of vertices of the morphing $(x_{i},y_{i}),(x_{i+1},y_{i+1})$ corresponds to two directed subsegments $\tau\subseteq\pi$ and $\nu\subseteq\sigma$. The matching of these two subsegments encodes a linear motion starting at the start points, and ending and the endpoints. See Figure 5.1.
@@
Given a morphing $\mathcalb{m}$, one can compute its width in linear time, as by convexity it is realized by the elevation of one of the vertices of $\mathcalb{m}$. Thus, a first step to trying to compute $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, is to compute the VE-Fréchet morphing $\mathcalb{m}_{\textsf{VE}}$ between $\pi$ and $\sigma$, and then directly monotonize it. If $\omega\left(\mathcalb{m}_{\textsf{VE}}\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$, then we have computed the Fréchet distance and so we are done. Otherwise, we will use refinement to narrow the interval $[\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$, as described below.
@@ 5.1.2 Monotonization via refinement
Let $\mathcalb{m}_{\textsf{VE}}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. By Observation 5.3, $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\in[\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$. Thus we have initial upper and lower bounds for $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and our approach is to repeatedly improve these bounds until they are equal, at which point the computed morphing realizes the Fréchet distance between the two curves.
@@
Indeed, create $\pi^{\prime}$ and $\sigma^{\prime}$ by introducing vertices along $\pi$ and $\sigma$ respectively, such that no edge has length exceeding $\varepsilon/2$. Recall that by Observation 4.9, $\mathcalb{m}_{\textsf{VE}}^{\prime}$ is already monotone over vertices, and thus $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}^{\prime}\right)$ does not change the edges from the curves that points in the morphing are mapped to. Thus the direct monotonizations of the VE morphings of $\pi^{\prime}$ and $\sigma^{\prime}$ can increase the distance by at most twice the maximum edge length.
@@
Let $r=\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and introduce a vertex in the interior of an edge of $\pi$ if this interior point is in distance exactly $r$ from some vertex of $\sigma$, and vice versa. Let $\pi^{\prime}$ and $\sigma^{\prime}$ be the resulting refined curves. Consider the grid $H$ of $R\left(\pi^{\prime},\sigma^{\prime}\right)$, and any edge $e$ of this grid. The value of the elevation function on $e$ is an interval: $\mathcalb{e}\left(e\right)=\left\{\mathcalb{e}\left(p\right)\;\middle|\;p\in e\right\}$, and this interval cannot contain $r$ in its interior (if it did, we would have broken it by introducing a vertex at the corresponding interior point). Thus any edge in $H$ whose elevation interval has a value above $r$, cannot contain any interior point with value below $r$, and thus such an edge is infeasible (in its interior) for the VE-Fréchet morphing $\mathcalb{m}_{\textsf{VE}}\left(\pi^{\prime},\sigma^{\prime}\right)$, as we know $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi^{\prime},\sigma^{\prime}\right)\leq r$. Namely, every edge of the free space is either completely feasible (for $r$), and or completely infeasible.
@@
It is straightforward to check if a given morphing $\mathcalb{m}$ is monotone. If it is not, then compute for each curve the portions where this morphing travels on them backward, and introduce a vertex in the middle of each backwards portion. Now, we recompute a morphing on the refined curve, Clearly this reduces the “back and forth” non-monotone motion on the two curves. One can naturally repeat this process several times till the non-monotonicity disappears, or becomes so small that it can be ignored. We refer to this as bisection refinement.
@@ 5.2 Computing the regular Fréchet distance
    First, compute the VE-Fréchet morphing. We now apply the bisection refinement, described in Remark 5.6, to the resulting morphing (recomputing the morphing using VE-Fréchet if needed). In practice, a few rounds of this bisection refinement makes these portions of the morphing involve leashes shorter than the bottleneck (even after direct monotonization), and we end up with the Fréchet distance.
@@
    Another (more systematic) approach to refinement is to consider all the vertices $V$ (say) of $\pi$ involved in a subpath of the morphing that is not monotone on an edge $e$ of $\sigma$. We refine $e$ along all the points that are intersections of bisectors of pair of points of $V$ with $e$. Since repeated application of this process converges to the refinement set of Lemma 5.5 (the set used in this lemma is too expensive to compute explicitly), this implies that sooner or later this yields the Fréchet distance (after direct monotonization). Essentially, this performs a search of the monotonicity events, see Section 1.2.2, that are relevant to the current morphing.

### v4

@@ 3.2.1 The discrete Fréchet distance
Let $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$ be two sequences of points in some normed space. Conceptually, we have two agents starting at $p_{1}$ and $q_{1}$ respectively. During an atomic time interval, one of them can jump forward to the next point in the sequence (one can allow both to jump forward in the same time, but we disallow it for the sake of simplicity of exposition). In the end, both agents have to arrive at $p_{n}$ and $q_{m}$, respectively, and our purpose is to minimize the maximum distance between them during this motion.
@@
Consider two sequences $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$, a discrete morphing is an $x/y$-monotone path $\tau=(1,1),\ldots,(n,m)$ in the grid graph defined over the set of points $U=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$ from $(1,1)$ to $(n,m)$. For a vertex $(i,j)\in U$, let $\mathcalb{h}\left(i,j\right)=\left\|{p_{i}q_{j}}\right\|$ be its associated weight. The path $\tau$ is restricted to vertical and horizontal edges of the grid. The width of $\tau$ is ${\omega}_{\pi,\sigma}\left(\tau\right)=\max_{(i,j)\in V\left(\tau\right)}\mathcalb{h}\left(i,j\right).$ The discrete Fréchet distance between $\pi$ and $\sigma$ is the minimum of the width of any morphing between the two sequences.
@@ 3.2.2 The retractable discrete Fréchet distance
Let $\tau=z_{1},\ldots,z_{n+m-1}\in U$ be a discrete morphing between two sequences $\pi=p_{1},\ldots,p_{n}$ and $\sigma=q_{1},\ldots,q_{m}$. Let $D(i,j)=\max_{k:i<k<j}\mathcalb{h}\left(z_{k}\right)$ be the inner width of the morphing $\tau(i,j)=z_{i},z_{i+1},\ldots,z_{j}$. For simplicity, assume all pairwise distances $\left\|{p_{i}q_{j}}\right\|$ are distinct. The unique point $\mathcalb{b}(i,j)$ realizing the minimum $D(i,j)$, overall possible grid monotone paths $\tau^{\prime}$ between $z_{i}$ and $z_{j}$, is the bottleneck between $i$ and $j$. The retractable discrete Fréchet morphing between $\pi$ and $\sigma$ is the unique morphing $\tau$, such that
@@
Although the weights here are defined on the vertices, it is easy enough to interpret them as being on the edges (by, for example, assigning a grid edge $z\rightarrow z^{\prime}$ the weight $\max(\mathcalb{h}\left(z\right),\mathcalb{h}\left(z^{\prime}\right))$. Plugging this (implicit) graph into Lemma 3.2 readily yields the following result.
@@
    The first running time bound of Lemma 3.5 is a worst-case bound, and for many realistic inputs, it is much smaller if the (discrete) Fréchet distance between them is small compared to the diameter of the two curves. See Figure 3.1 for an example.
@@ 4.1 Definition and basic algorithm
Let $\mathcalb{p}_{i}=\overrightarrow{p_{i}p_{i+1}}$ and $\mathcalb{q}_{j}=\overrightarrow{q_{j}q_{j+1}}$. Let $x_{i}=\left\|{\pi[p_{1},p_{i}]}\right\|$ and $y_{j}=\left\|{\sigma[q_{1},q_{j}]}\right\|$. The grid cell $C_{i,j}$ is the subrectangle $x_{i},x_{i}+\left\|{p_{i}p_{i+1}}\right\|]\times[y_{j},y_{j}+\left\|{q_{j}q_{j+1}}\right\|]$. The elevation function for any point $(x,y)$ inside $C_{i,j}$ is given by
@@
We emphasize that the running time bound in this theorem is pessimistic, and in many cases, the algorithm is significantly faster. For an example of the output of the algorithm, see Figure 4.2.
Figure 4.2: The VE-Fréchet morphing for the two curves from Figure 1.3. Note that the solution is “slightly” not $x/y$-monotone. For animation of this morphing, follow the link.
@@ 4.1.1 Reachable complexity
A similar notion (relative free space complexity) was used by Driemel et al. [DHW12]. If $r=\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)$, then the algorithm for computing the VE morphing only explores vertices in the $r$-reachable region, which readily implies the following.
@@
The above implies that the worst case, for the above algorithm, is when $N=nm$. This happens, for example, (somewhat counterintuitively) when $\pi$ and $\sigma$ are far from each other, and their VE-Fréchet distance is realized by the distance of their endpoints (say, at the start of the morphing). This somewhat absurd situation can happen for real-world inputs – follow this link for an example. Furthermore, this quadratic behavior is probably unavoidable, as computing the VE-Fréchet distance probably requires quadratic time in the worst case. We offer some approaches to avoid this for real-world input later on.
@@ 4.2 Basic properties of the VE Fréchet morphing/distance
Specifically, let $\mathcalb{m}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. If we project $\mathcalb{m}$ on the $x$-axis, denoted by $x(\mathcalb{m})$, we get a path from 0 to $\left\|{\pi}\right\|$ on the real line. Note, that this path is not necessarily monotone, but importantly, for any vertex $v\in\pi$ it is monotone. Specifically, let $\alpha$ be the distance of $v$ along $\pi$ from its start. Then, the path $x(\mathcalb{m})$ crosses from one side of $\alpha$ to the other side, exactly once (assuming $v$ is an internal vertex). To see why this is true, observe that any morphing must cross the vertical line $x=\alpha$. This line has vertices along it, and observe that in $G_{\textsf{VE}}\left(\pi,\sigma\right)$ it is only possible to cross this line by using one of these vertices, as no edge crosses from one side of this line to the other (i.e. the vertices on this line are a separator). However, vertices along this line have outgoing edges only to vertices that are to the right (or remain on the same $x$, but these target vertices are on grid edges adjacent to the right of this line).
@@ 5.1 Direct monotonization
If the VE-Fréchet morphing is $x/y$-monotone, then it realizes the (regular) Fréchet distance between the curves. Indeed, if the morphing is monotone, then it is a valid morphing considered by Definition 1.4 for the Fréchet distance, and thus $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)\geq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\geq\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right)$, where the latter inequality is from Lemma 4.8.
@@
A VE-Fréchet morphing $\mathcalb{m}$ in a free space diagram $R=R\left(\pi,\sigma\right)$ of two curves $\pi$ and $\sigma$ with $n$ and $m$ vertices, respectively, is a polygonal path with $t=n+m+O(1)$ vertices. Thus, a morphing can be represented as $\mathcalb{m}=(x_{1},y_{1}),\ldots,(x_{t},y_{t})$. A consecutive pair of vertices of the morphing $(x_{i},y_{i}),(x_{i+1},y_{i+1})$ corresponds to two directed subsegments $\tau\subseteq\pi$ and $\nu\subseteq\sigma$. The matching of these two subsegments encodes a linear motion starting at the start points and ending at the endpoints. See Figure 5.1.
@@
Given a morphing $\mathcalb{m}$, one can compute its width in linear time, as by convexity it is realized by the elevation of one of the vertices of $\mathcalb{m}$. Thus, a first step to trying to compute $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$ is to compute the VE-Fréchet morphing $\mathcalb{m}_{\textsf{VE}}$ between $\pi$ and $\sigma$, and then directly monotonize it. If $\omega\left(\mathcalb{m}_{\textsf{VE}}\right)=\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)$, then we have computed the Fréchet distance, and so we are done. Otherwise, we will use refinement to narrow the interval $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$, as described below.
@@ 5.1.2 Monotonization via refinement
Let $\mathcalb{m}_{\textsf{VE}}$ be the VE-Fréchet morphing between $\pi$ and $\sigma$. By Observation 5.3, $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)\in[\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi,\sigma\right),\omega\left(\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}\right)\right)]$. Thus, we have initial upper and lower bounds for $\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$. The idea is to improve, repeatedly, these bounds until they are equal. At this point, the computed morphing realizes the Fréchet distance between the two curves.
@@
Indeed, create $\pi^{\prime}$ and $\sigma^{\prime}$ by introducing vertices along $\pi$ and $\sigma$ respectively, such that no edge has length exceeding $\varepsilon/2$. Recall that by Observation 4.9, $\mathcalb{m}_{\textsf{VE}}^{\prime}$ is already monotone over vertices, and thus $\mathrm{mono}\left(\mathcalb{m}_{\textsf{VE}}^{\prime}\right)$ does not change the edges from the curves that points in the morphing are mapped to. Thus, the direct monotonizations of the VE morphings of $\pi^{\prime}$ and $\sigma^{\prime}$ can increase the distance by at most twice the maximum edge length.
@@
Let $r=\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, and introduce a vertex in the interior of an edge of $\pi$ if this interior point is in distance exactly $r$ from some vertex of $\sigma$, and vice versa. Let $\pi^{\prime}$ and $\sigma^{\prime}$ be the resulting refined curves. Consider the grid $H$ of $R\left(\pi^{\prime},\sigma^{\prime}\right)$, and any edge $e$ of this grid. The value of the elevation function on $e$ is an interval: $\mathcalb{e}\left(e\right)=\left\{\mathcalb{e}\left(p\right)\;\middle|\;p\in e\right\}$, and this interval cannot contain $r$ in its interior (if it did, we would have broken it by introducing a vertex at the corresponding interior point). Thus, any edge in $H$ whose elevation interval has a value above $r$ cannot contain any interior point with value below $r$. Thus, such an edge is infeasible (in its interior) for the VE-Fréchet morphing $\mathcalb{m}_{\textsf{VE}}\left(\pi^{\prime},\sigma^{\prime}\right)$, as we know $\mathsf{d}^{ve}_{\mathcal{F}}\left(\pi^{\prime},\sigma^{\prime}\right)\leq r$. Namely, every edge of the free space is either entirely feasible (for $r$) or infeasible.
@@
It is straightforward to check if a given morphing $\mathcalb{m}$ is monotone. If it is not, then compute for each curve the portions where this morphing travels on them backward, and introduce a vertex in the middle of each backward portion. Now, we recompute a morphing on the refined curve. Clearly, this reduces the “back and forth” non-monotone motion on the two curves. One can naturally repeat this process several times till the non-monotonicity disappears or becomes so small that it can be ignored. We refer to this as bisection refinement.
@@ 5.2 Computing the regular Fréchet distance
    First, compute the VE-Fréchet morphing. We now apply the bisection refinement, described in Remark 5.6, to the resulting morphing (recomputing the morphing using VE-Fréchet if needed). In practice, a few rounds of this bisection refinement make these portions of the morphing involve leashes shorter than the bottleneck (even after direct monotonization), and we end up with the Fréchet distance.
@@
    Another (more systematic) approach to refinement is to consider all the vertices $V$ (say) of $\pi$ involved in a subpath of the morphing that is not monotone on an edge $e$ of $\sigma$. We refine $e$ along all the points that are intersections of bisectors of pairs of points of $V$ with $e$. Since repeated application of this process converges to the refinement set of Lemma 5.5 (the set used in this lemma is too expensive to compute explicitly), this implies that sooner or later this yields the Fréchet distance (after direct monotonization). Essentially, this performs a search of the monotonicity events, see Section 1.2.2, that are relevant to the current morphing.